Coulomb's law for the magnitude of the force F between two particles with charges Q and Q′ separated by a distance d is |F|=K|QQ′|d2, where K=14πϵ0, and ϵ0=8.854×10−12C2/(N⋅m2) is the permittivity of free space. Consider two point charges located on the x axis: one charge, q1 = -16.0 nC , is located at x1 = -1.675 m ; the second charge, q2 = 37.0 nC , is at the origin (x=0.0000). Part A What is the net force exerted by these two charges on a third charge q3 = 50.0 nC placed between q1 and q2 at x3 = -1.120 m ? Your answer may be positive or negative, depending on the direction of the force. Express your answer numerically in newtons to three significant figures.
Answer #1

9x107 IFI= k QQ 3 4* Lo q = -16oone (2=-1-675m) q2 = 37.one (22=0) Part A (x2 = -1120) V = 50.one <- 4:175 -1.1200 9, 93 > V